# DiffInk: Glyph- and Style-Aware Latent Diffusion Transformer for Text to Online Handwriting Generation

### v2

@@ Abstract
Deep generative models have advanced text-to-online handwriting generation (TOHG), which aims to synthesize realistic pen trajectories conditioned on textual input and style references. However, most existing methods still primarily focus on character- or word-level generation, resulting in inefficiency and a lack of holistic structural modeling when applied to full text lines. To address these issues, we propose DiffInk, the first latent diffusion Transformer framework for full-line handwriting generation. We first introduce InkVAE, a novel sequential variational autoencoder enhanced with two complementary latent-space regularization losses: (1) an OCR-based loss enforcing glyph-level accuracy, and (2) a style-classification loss preserving writing style. This dual regularization yields a semantically structured latent space where character content and writer styles are effectively disentangled. We then introduce InkDiT, a novel latent diffusion Transformer that integrates target text and reference styles to generate coherent pen trajectories. Experimental results demonstrate that DiffInk outperforms existing state-of-the-art methods in both glyph accuracy and style fidelity, while significantly improving generation efficiency.
@@ 1 Introduction
In recent years, TOHG has attracted growing research interest. Early methods (Graves, 2013; Li et al., 2024) adopt Auto-Regressive (AR) modeling to generate trajectory points sequentially. Subsequent approaches attempt to control style through content-style disentanglement, utilizing either 1) coarse-grained global style representations such as writer IDs (Luo et al., 2022) or 2) fine-grained local style representations such as character components (Liu et al., 2024; Dai et al., 2023). Notably, OLHWG (Ren et al., 2025) adopts a novel diffusion-based framework to generate isolated characters and arranges them using external layout prediction modules, achieving state-of-the-art performance. However, these methods all fundamentally operate at the character or word level, lacking the ability to capture long-range dependencies and overall text-line structure.
@@
Our main contributions can be summarized as follows: (1) We propose DiffInk, the first latent diffusion framework for end-to-end full-line handwriting generation, capable of generating glyph-accurate and style-consistent pen trajectories. (2) We introduce InkVAE, a sequential VAE enhanced with two task-relevant lightweight regularization losses (OCR-based loss and style-classification loss) to disentangle content and style in the latent space, enabling structured representation learning. (3) We propose a novel InkDiT that jointly conditions on target text and reference style, generating realistic handwriting trajectories via an iterative denoising process. (4) Experiments on the CASIA-OLHWDB 2.0–2.2 benchmark demonstrate that DiffInk significantly outperforms current state-of-the-art methods in full-line generation quality, while also offering improved efficiency.
@@ 2.3 Diffusion Generative Models
Diffusion models (Ho et al., 2020) have emerged as powerful tools in generative modeling, demonstrating impressive performance across images (Rombach et al., 2022), speech (Jiang et al., 2025), and vision-language tasks (Luo et al., 2023). To reduce computational costs and enhance semantic control, many recent works (Zhang et al., 2023; Saharia et al., 2022; Ramesh et al., 2022; Labs, 2024) adopt latent diffusion models, where the denoising process is conducted in a learned latent space. Recently, Diffusion Transformers (DiT) (Peebles and Xie, 2023) have replaced traditional U-Net (Ronneberger et al., 2015) backbones with Transformer-based (Vaswani et al., 2017) architectures, enabling stronger global conditioning in image synthesis tasks. Variational Autoencoders (VAEs) (Van Den Oord et al., 2017; Baevski et al., 2020) are commonly used to construct such spaces, offering compact and structured representations.
@@ 4.3 Main Results
As there is a lack of established baselines for end-to-end text-line generation, we adopt several state-of-the-art character generation methods and extend them with a layout modeling module to enable line-level comparison. These baselines include Drawing (Zhang et al., 2017), DeepImitator (Zhao et al., 2020), WLU (Tang and Lian, 2021), and SDT (Dai et al., 2023), which generate high-fidelity handwritten characters under an autoregressive paradigm, as well as OLHWG (Ren et al., 2025), which synthesizes isolated characters using a diffusion-based approach. To support text-line level comparison, we adopt the layout prediction module from OLHWG—currently the state-of-the-art—to compose full lines from the character-level outputs of all five methods. In the OLHWG method, the layout prediction module infers the subsequent layout based on the layouts of a fixed number of preceding characters (originally set to 10) from the ground-truth text line. To accommodate text lines of varying lengths, we instead use the layouts of the first 30% of characters from each text line as input for layout prediction (max 10). To ensure fairness, all methods are retrained on the same dataset with consistent preprocessing and default configurations.
@@
As shown in Table 1, DiffInk (ours) achieves superior performance across all quantitative metrics. For content fidelity, the OCR-based metrics AR and CR both exceed 94%, outperforming the latest state-of-the-art method OLHWG by an average margin of 3 percentage points, and substantially surpassing earlier approaches such as Drawing, DeepImitator, WLU and SDT—with a margin of 14.44%. For style consistency, DiffInk achieves a style score of 77.38 at the text-line level—the highest among all methods—surpassing the contrastive-learning–based OLHWG by 30 percentage points, the dual-branch style-loss–based SDT by 27 points, and baselines with simple style modeling by an average of 50 points. Moreover, style evaluation further reveals that layout stitching introduces anomalies that can be readily detected by the style classifier. In terms of structural fidelity, DiffInk attains substantially lower normalized DTW distances, indicating more consistent alignment with target trajectories. In contrast, autoregressive methods such as WLU and SDT suffer from cumulative error propagation, which severely degrades text-line level performance. Finally, DiffInk demonstrates strong generation efficiency, producing 58.82 characters per second—17× faster than SDT, over 800× faster than OLHWG, and more than 2× faster than WLU—thanks to the efficient text-line modeling of InkVAE.
@@ 5 Conclusion
The reproduction procedure consists of the following steps: (1) Prepare the required text-line data as described in Section A; (2) Implement the DiffInk framework, including InkVAE and InkDiT, following the methodology (Sections 3.3 and 3.4) and the framework details in Section B; (3) Train InkVAE end-to-end according to the implementation details in Section 4.2; (4) Train InkDiT following the implementation details in Section 4.2 and the discussion in Section D. For evaluation and comparison with other methods, please refer to Sections 4.3 and B.4.
@@ C.2 More Comparisons with existing Methods
We provide additional comparisons with existing approaches, including Drawing (Zhang et al., 2017), Deepiminator (Zhao et al., 2020)WLU (Tang and Lian, 2021), SDT (Dai et al., 2023), and OLHWG (Ren et al., 2025). As illustrated in Figure 12 and Figure 13, our method delivers superior text-line generation quality over a broader spectrum of handwriting styles.
@@ D.6 Limitations
While our model establishes state-of-the-art performance on the notoriously difficult task of Chinese handwritten text-line generation, direct application to other languages (e.g., Latin-based scripts) still requires retraining. Exploring how to develop a unified multilingual generation model with only minor modifications will therefore be one of our key directions for future work.
@@ Appendix E Use of Large Language Models (LLMs)
LLMs, such as ChatGPT (OpenAI, 2023) and DeepSeek (AI, 2024), were employed solely for language polishing and minor writing refinements in this paper. They were not used for generating ideas, designing methods, conducting experiments, analyzing results, or creating figures. All scientific contributions, including the conception of the research problem, model development, and experimental validation, were carried out independently by the authors.

### v3

@@ Abstract
Deep generative models have advanced text-to-online handwriting generation (TOHG), which aims to synthesize realistic pen trajectories conditioned on textual input and style references. However, most existing methods still primarily focus on character- or word-level generation, resulting in inefficiency and a lack of holistic structural modeling when applied to full text lines. To address these issues, we propose DiffInk, the first latent diffusion Transformer framework for full-line handwriting generation. We first introduce InkVAE, a novel sequential variational autoencoder enhanced with two complementary latent-space regularization losses: (1) an OCR-based loss enforcing glyph-level accuracy, and (2) a style-classification loss preserving writing style. This dual regularization yields a semantically structured latent space where character content and writer styles are effectively disentangled. We then introduce InkDiT, a novel latent diffusion Transformer that integrates target text and reference styles to generate coherent pen trajectories. Experimental results demonstrate that DiffInk outperforms existing state-of-the-art (SOTA) methods in both glyph accuracy and style fidelity, while significantly improving generation efficiency.
@@ 1 Introduction
In recent years, TOHG has attracted growing research interest. Early methods (Graves, 2013; Li et al., 2024) adopt Auto-Regressive (AR) modeling to generate trajectory points sequentially. Subsequent approaches attempt to control style through content-style disentanglement, utilizing either 1) coarse-grained global style representations such as writer IDs (Luo et al., 2022) or 2) fine-grained local style representations such as character components (Liu et al., 2024; Dai et al., 2023). Notably, OLHWG (Ren et al., 2025) adopts a novel diffusion-based framework to generate isolated characters and arranges them using external layout prediction modules, achieving SOTA performance. However, these methods all fundamentally operate at the character or word level, lacking the ability to capture long-range dependencies and overall text-line structure.
@@
Our main contributions can be summarized as follows: (1) We propose DiffInk, the first latent diffusion framework for end-to-end full-line handwriting generation, capable of generating glyph-accurate and style-consistent pen trajectories. (2) We introduce InkVAE, a sequential VAE enhanced with two task-relevant lightweight regularization losses (OCR-based loss and style-classification loss) to disentangle content and style in the latent space, enabling structured representation learning. (3) We propose a novel InkDiT that jointly conditions on target text and reference style, generating realistic handwriting trajectories via an iterative denoising process. (4) Experiments on the CASIA-OLHWDB 2.0–2.2 benchmark demonstrate that DiffInk significantly outperforms current SOTA methods in full-line generation quality, while also offering improved efficiency.
@@ 2.3 Diffusion Generative Models
Diffusion models (Ho et al., 2020) have emerged as powerful tools in generative modeling, demonstrating impressive performance across images (Rombach et al., 2022), speech (Lee et al., 2025), and vision-language tasks (Luo et al., 2023). To reduce computational costs and enhance semantic control, many recent works (Zhang et al., 2023; Saharia et al., 2022; Ramesh et al., 2022; Labs, 2024) adopt latent diffusion models, where the denoising process is conducted in a learned latent space. Recently, Diffusion Transformers (DiT) (Peebles and Xie, 2023) have replaced traditional U-Net (Ronneberger et al., 2015) backbones with Transformer-based (Vaswani et al., 2017) architectures, enabling stronger global conditioning in image synthesis tasks. Variational Autoencoders (VAEs) (van den Oord et al., 2017; Baevski et al., 2020) are commonly used to construct such spaces, offering compact and structured representations.
@@ 4.3 Main Results
As there is a lack of established baselines for end-to-end text-line generation, we adopt several SOTA character generation methods and extend them with a layout modeling module to enable line-level comparison. These baselines include Drawing (Zhang et al., 2017), DeepImitator (Zhao et al., 2020), WLU (Tang and Lian, 2021), and SDT (Dai et al., 2023), which generate high-fidelity handwritten characters under an autoregressive paradigm, as well as OLHWG (Ren et al., 2025), which synthesizes isolated characters using a diffusion-based approach. To support text-line level comparison, we adopt the layout prediction module from OLHWG—currently the SOTA—to compose full lines from the character-level outputs of all five methods. In the OLHWG method, the layout prediction module infers the subsequent layout based on the layouts of a fixed number of preceding characters (originally set to 10) from the ground-truth text line. To accommodate text lines of varying lengths, we instead use the layouts of the first 30% of characters from each text line as input for layout prediction (max 10). To ensure fairness, all methods are retrained on the same dataset with consistent preprocessing and default configurations.
@@
As shown in Table 1, DiffInk (ours) achieves superior performance across all quantitative metrics. For content fidelity, the OCR-based metrics AR and CR both exceed 94%, outperforming the latest SOTA method OLHWG by an average margin of 3 percentage points, and substantially surpassing earlier approaches such as Drawing, DeepImitator, WLU and SDT—with a margin of 14.44%. For style consistency, DiffInk achieves a style score of 77.38 at the text-line level—the highest among all methods—surpassing the contrastive-learning–based OLHWG by 30 percentage points, the dual-branch style-loss–based SDT by 27 points, and baselines with simple style modeling by an average of 50 points. Moreover, style evaluation further reveals that layout stitching introduces anomalies that can be readily detected by the style classifier. In terms of structural fidelity, DiffInk attains substantially lower normalized DTW distances, indicating more consistent alignment with target trajectories. In contrast, autoregressive methods such as WLU and SDT suffer from cumulative error propagation, which severely degrades text-line level performance. Finally, DiffInk demonstrates strong generation efficiency, producing 58.82 characters per second—17× faster than SDT, over 800× faster than OLHWG, and more than 2× faster than WLU—thanks to the efficient text-line modeling of InkVAE.
@@ 5 Conclusion
Acknowledgement
This research is supported in part by the National Natural Science Foundation of China (Grant No.:62476093).
@@ Acknowledgement
Reproduction consists of dataset preparation (Sections 4.1 and Appendix A), the design of DiffInk (InkVAE and InkDiT) as described in Sections 3.3 and 3.4, implementation and training detailed in Appendix B and Section 4.2, and evaluation following Section 4.3 and Appendix B.4.
@@ C.2 More Comparisons with existing Methods
We provide additional comparisons with existing approaches, including Drawing (Zhang et al., 2017), Deepiminator (Zhao et al., 2020), WLU (Tang and Lian, 2021), SDT (Dai et al., 2023), and OLHWG (Ren et al., 2025). As illustrated in Figure 12 and Figure 13, our method delivers superior text-line generation quality over a broader spectrum of handwriting styles.
@@ D.6 Limitations
While our model establishes SOTA performance on the notoriously difficult task of Chinese handwritten text-line generation, direct application to other languages (e.g., Latin-based scripts) still requires retraining. Exploring how to develop a unified multilingual generation model with only minor modifications will therefore be one of our key directions for future work.
@@ Appendix E Use of Large Language Models (LLMs)
LLMs, such as ChatGPT (OpenAI, 2023) and DeepSeek (DeepSeek-AI, 2024), were employed solely for language polishing and minor writing refinements in this paper. They were not used for generating ideas, designing methods, conducting experiments, analyzing results, or creating figures. All scientific contributions, including the conception of the research problem, model development, and experimental validation, were carried out independently by the authors.